Sol Summary – February 13th

Crew 221 Sol Summary Report 13Feb-2020


Summary Title: Real Mars Experience

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Today the crew woke up and had pancakes prepared by the crew commander and with the olive oil we got from the last supply mission from Earth. Finally, they do not stick to the pan anymore and were delicious!

In the morning we were under the gun to get the robot ready as there were hookups with the software. All bugs were solved by 13:00, leaving one hour before the EVA.

The girls EVA team, containing the crew commander M.Grulich, crew journalist H. Kalucha and crew GreenHab Officer J. Todd who mainly worked on the robot, got ready for the EVA. They decided to test the robot just in front of the hab door to make sure it was working. It did but the wheels came loose, and it had to be fixed in the field. This was a real challenge, requesting several tools from the team support inside the hab.

Once all issues were fixed the EVA team headed for the rover but realized that Spirit was not plugged in. After radio communication, it was decided to leave one team member at the hab and go with permission with two crew members on Curiosity. Crew commander and crew GreenHab officer drove to Steambed Connector and walked to Candor Chasma and tested the robot successfully in a more challenging environment.

Once back the team had a debrief discussing the communication and confusion of information that lead to one person not joining the EVA mission and realized that any piece of information from mission support should and shall in the future be reviewed and read in more detail, as we cannot rely on mission support on Mars.

The evening was spent with a lovely salad prepared by crew HSO Rawan Alshammari and will be concluded with a games night.

Anomalies in work: None

Update mouse in hab: no new mice caught

Weather: sunny light winds

Crew Physical Status: Crew some sore throat, but seems to improve

Reports to be filed:

Sol Summary
Photo Report
Operations Report
Journalist Report
GreenHab Report
EVA Request
EVA Report 13

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site